### v2

# Improving Sequential Market Coordination via Value-oriented Renewable Energy Forecasting

Abstract
Large penetration of renewable energy sources (RESs) brings huge uncertainty into the electricity markets. The current deterministic clearing approach in the day-ahead (DA) market, where RESs participate based on expected production, has been criticized for causing a lack of coordination between the DA and real-time (RT) markets, leading to high overall operating costs. Previous works indicate that improving day-ahead RES entering quantities can significantly mitigate the drawbacks of deterministic clearing. In this work, we propose using a trained forecasting model, referred to as value-oriented forecasting, to determine RES Improved Entering Quantities (RIEQ) more efficiently during the operational phase. Unlike traditional models that minimize statistical forecasting errors, our approach trains model parameters to minimize the expected overall operating costs across both DA and RT markets. We derive the exact form of the loss function used for training, which becomes piecewise linear when market clearing is modeled by linear programs. Additionally, we provide the analytical gradient of the loss function with respect to the forecast, enabling an efficient training strategy. Numerical studies demonstrate that our forecasts significantly reduce overall operating costs for deterministic market clearing compared to conventional forecasts based on expected RES production.
Keywords: Energy forecasting, Loss function, Forecast value, Market clearing, Decision-focused learning
I Introduction
The current short-term electricity markets are organized in a sequence of trading floors, i.e., day-ahead (DA) and real-time (RT) markets [1]. A DA market is cleared 12-36 hours before the actual operation. A RT market runs close to the delivery time and addresses any imbalance from the DA schedules. They are initially designed for controllable fossil-fueled generators in the view of traditional power system operation. However, the increasing share of renewable energy sources (RESs) (up to 30% of global electricity generation in 2022 [2]) exposes the electricity markets to significant uncertainty and therefore raises concerns to the market operation [1].
A significant challenge with sequential deterministic market clearing arises from its limited ability to coordinate the two markets under large RES uncertainty. Currently, the DA market is cleared without accounting for its impact on RT market clearing. As a result, DA and RT markets operate without coordination, leading to high overall operating costs [3]. Consequently, stochastic market clearing has been proposed, which informs the DA market of the operating cost in the RT market [4]. While stochastic market clearing can improve overall economic efficiency, it struggles to simultaneously maintain key market properties, such as revenue adequacy and cost recovery [5]. To address this, efforts have been made to preserve these desirable properties within stochastic market clearing. For instance, [6] guarantees cost recovery and revenue adequacy both per scenario and in expectation, though this comes at the cost of reduced market efficiency.
Previous studies have shown that combining DA deterministic market clearing with RES improved entering quantities (RIEQ) enhances coordination between DA and RT markets, achieving performance comparable to stochastic clearing [7, 8]. The key idea is to tactically determine RIEQ—used as input parameters in DA clearing—by considering their impact on future RT market operations. For instance, studies [7, 8] optimize RIEQ by solving a bilevel program on a case-by-case basis during the operational phase. Similarly, [9] applies these methods to predict reserve requirements, facilitating coordination of reserve and energy markets. While these approaches hold promise for improving market coordination, they introduce computational challenges in determining RIEQ.
Building on this foundation, the goal of this work is not to replace the existing RES entering quantities determined by RES producer offerings in the DA market. Instead, we aim to explore an important question: What if the RES entering quantities in the DA market were improved as RIEQ by a “properly” trained forecasting model, specifically optimized to minimize overall system operating costs? If so, could the sequential markets be better coordinated? Here, the overall system operating cost in both the DA and RT markets is used as the quantitative metric to evaluate market coordination.
With the trained forecasting model for determining RIEQ, the computational burden at the operational phase [7, 8] can be avoided. However, aligning the training objective of a forecasting model with the operational goal is a significant challenge, falling within the realm of value-oriented forecasting [10, 11, 12]. Several research threads have emerged to address this challenge, encompassing integrated optimization, differentiable programming, and the loss function design.
For the first thread of research on integrated optimization, forecasting model parameters are optimized concurrently with decision variables [11, 13]. In the same vein, [12] introduces a bilevel program where DA market clearing forms the lower level with the RIEQ prediction as a parameter, while the upper level optimizes both model parameters and RT decisions. This method requires a linear forecasting model to make the program solvable by commercial optimization solvers, which may limit the performance. The second thread of research, on differentiable programming, enables the use of more advanced forecasting models, such as neural networks, by deriving the gradient of decision solution w.r.t. the forecast [14, 15]. However, obtaining such a gradient involves repeatedly solving the inverse of a large-size matrix resulting from the Karush–Kuhn–Tucker conditions, making it computationally expensive. The third approach, and the focus of our work, centers on designing value-oriented loss functions. The existing ones, such as pinball loss [16] or Smart “Predict, then Optimize” (SPO) loss [17], are primarily designed for single-stage stochastic programs (e.g., the one without considering flexible unit redispatch in the RT market). Thus, how to develop a value-oriented loss function for sequential market clearing remains an open question.
To address the challenge of training the RIEQ forecasting model, we propose a tailored loss function that uses the RIEQ forecast as the input and the overall DA and RT system operating costs as the output. The forecasting model parameter estimation is formulated as a bilevel program: the upper level optimizes forecasting model parameters, while the lower level solves the DA and RT market clearing problems using the RIEQ forecast as the input parameter. To link operating cost with the RIEQ forecast, we leverage the lower-level dual problems and replace the upper-level objective with dual objectives. The reformulated upper-level objective is then transformed into an analytical function of the forecast, serving as the loss function for training. This requires the derivation of functions that connect the forecast with primal and dual solutions. Specifically, the dual solutions are linked to forecasts through the dual problems of DA and RT market clearing, while the primal solutions are derived from the active constraints of their respective primal problems. Substituting these functions into the upper-level objective yields a value-oriented loss function for training. Our main contributions are,
1) From a market perspective, the proposed approach maintains deterministic DA market clearing while improving sequential market coordination with RIEQ forecasts.
2) From a theoretical perspective, we derive a value-oriented loss function that aligns the forecasting model’s training objective with operational value, minimizing overall system operating costs in both DA and RT markets.
3) From a practitioner’s perspective, the analytical loss function allows the analytical derivation of the gradient. Leveraging the structure of the gradient, we propose a computationally efficient training approach.
Although prior works [18, 19] design forecasting products using value-oriented loss functions, these approaches are not directly applicable to electricity market settings. First, prior works cannot address the participation of entities other than RES in both DA and RT markets, which is typical in electricity markets. As a result, the network and the redispatch of flexible resources cannot be handled. Second, the generalized market setting introduces new technical challenges in obtaining the value-oriented loss function. The reformulated upper-level objective, derived from dual problems, depends on both the forecast and the primal and dual solutions from the lower level. This requires us to identify the functions between those solutions and the forecast.
The remaining parts of this paper are organized as follows. The preliminaries regarding the sequential market clearing are given in Section II. Section III formulates a bilevel program for forecasting model parameter estimation. Section IV derives the loss function for value-oriented forecasting and the training process is presented in Section V. Results are discussed and evaluated in Section VI, followed by the conclusions.
Notations: The notation $\bm{X}[\mathcal{J}]$ signifies the sub-matrix comprised of rows from the matrix $\bm{X}$ whose indices are included in the index set $\mathcal{J}$. Considering the column vectors $\bm{x}_{1}$ and $\bm{x}_{2}$, the expression $\bm{x}=[\bm{x}_{1};\bm{x}_{2}]$ indicates the vertical concatenation of $\bm{x}_{1}$ and $\bm{x}_{2}$, resulting in the formation of $\bm{x}$. The operator $\Pi_{\mathcal{J}}$ is used to extract a segment from the vector, such as retrieving $\bm{x}_{1}$ from $\bm{x}$.
II Preliminaries
The framework and mathematical formulation of sequential market clearing are introduced in subsection II-A, and the reformulation is presented in subsection II-B.
Figure 1: An illustration of forecasting and the sequential market clearing where $\hat{\bm{y}}_{d,\tau}$ and $\bm{y}_{d,\tau}$ are forecasts and realization, and $\bm{w}_{d,\tau}^{*}$ is the RES schedule.
II-A Sequential Market Clearing
We consider the sequential clearing of DA and RT markets [1], as illustrated in Fig. 1. The DA market is cleared at time $t$ on day $d-1$, with an advance of $k$ hours in time to the next day $d$, and covers energy transactions on day $d$, typically on an hourly basis. RES production is uncertain during the DA market, where the RES forecast acts as the RIEQ. Due to inevitable forecasting errors, the energy imbalance caused by forecasting errors needs to be settled in the RT market. Concretely, in the DA market, the operator determines the schedules of generators and RES to satisfy inelastic demand. The generation and RES schedule for each time-slot $\tau,\forall\tau=1,...,24$ on the next day $d$ are denoted as $\bm{p}_{d,\tau}$ and $\bm{w}_{d,\tau}$, respectively. The DA market clearing is,
$\mathop{\min}_{\bm{x}_{d}}$ $\sum_{\tau=1}^{T}\bm{\rho}^{\top}\bm{p}_{d,\tau}$ (1a)
s.t. $\bm{1}^{\top}(\bm{p}_{d,\tau}+\bm{w}_{d,\tau})=\bm{1}^{\top}\bm{l}_{d,\tau},\forall\tau=1,...,T$ (1b)
$-\overline{\bm{f}}\leq\bm{H}(\bm{p}_{d,\tau}+\bm{w}_{d,\tau}-\bm{l}_{d,\tau})\leq\overline{\bm{f}},\forall\tau=1,...,T$ (1c)
$\underline{\bm{p}}\leq\bm{p}_{d,\tau}\leq\overline{\bm{p}},\forall\tau=1,...,T$ (1d)
$\underline{\bm{r}}\leq\bm{p}_{d,\tau}-\bm{p}_{d,\tau-1}\leq\overline{\bm{r}},\forall\tau=2,...,T$ (1e)
$0\leq\bm{w}_{d,\tau}\leq\hat{\bm{y}}_{d,\tau},\forall\tau=1,...,T,$ (1f)
where $\bm{x}_{d}=[\bm{x}_{d,\tau}]_{\tau=1}^{T}=[\bm{p}_{d,\tau}\in\mathbb{R}^{N};\bm{w}_{d,\tau}\in\mathbb{R}^{N}]_{\tau=1}^{T}$ is the collection of DA decision variables, and $N$ is the number of nodes in the system. The optimal solution is denoted as $\bm{x}_{d}^{*}=[\bm{x}_{d,\tau}^{*}]_{\tau=1}^{T}=[\bm{p}_{d,\tau}^{*};\bm{w}_{d,\tau}^{*}]_{\tau=1}^{T}$. $\bm{\rho}\in\mathbb{R}^{N}$ is the marginal cost vector of traditional generators. RES enters the market with zero marginal cost. Each element in the vector $\bm{p}_{d,\tau}$ represents the power generated by a traditional generator unit, whose marginal cost is in the corresponding element in $\bm{\rho}$. Here, we assume there is no power loss on lines, and include a DC representation of the network. The equality constraint (1b) enforces the power balance conditions. For simplification and considering the high accuracy of demand forecast, the demand $\bm{l}_{d,\tau}\in\mathbb{R}^{N}$ is considered to be known with certainty. The inequality constraints (1c) restrict the scheduled power flow within the line flow limits. $\bm{H}$ in (1c) is the Power Transfer Distribution Factors [20] mapping the nodal power injection to the power flow on lines. (1d) and (1e) are the output power and ramping limits of the traditional generators. (1f) limits the DA schedule of RES up to the forecast $\hat{\bm{y}}_{d,\tau}$ which represents a single-value estimate of the RES production $Y_{d,\tau}$. $Y_{d,\tau}$ is a random variable since the RES production is unknown in DA market clearing.
Remark 1.
In line with European practices [12, 21], we do not incorporate binary decisions regarding unit commitment (UC) in (1). However, we note that UC is a requisite consideration in the U.S. markets. To analyze the market behavior, the relaxed UC problem, where the binary commitment decisions are substituted with continuous ones, is widely used [3, 8]. The relaxed UC is a linear program. A common practice is to train the forecasting model using the relaxed UC formulation [8]. During the operational phase, the generated forecasts are then used as inputs to the UC problem.
Since the RES production is uncertain in DA, the DA schedules are to be adjusted at each time-slot $\tau,\forall\tau=1,...,T$ in RT on day $d$, after the RES realization $\bm{y}_{d,\tau}$ is observed. The RT market deals with the imbalance $\bm{y}_{d,\tau}-\bm{w}_{d,\tau}^{*}$ caused by RES, with a minimized imbalance cost. Additionally, the RT market clearing at time-slot $\tau,\forall\tau=2,...,T$ is influenced not only by the DA clearing outcomes but also by the extent of power adjustments made in the preceding time-slot. This is due to the ramping constraints that interconnect adjacent time slots. Because the market clearing is conducted separately for each day, the RT market clearing at time-slot $\tau=1$ on day $d$ remains unaffected by any adjustments made at time-slot $\tau=T$ on the previous day $d-1$. In the following, we firstly give the mathematical formulation of the RT clearing at $\tau=1$,
$\mathop{\min}_{\bm{z}_{d,\tau}}$ $\ \bm{\rho}_{+}^{\top}\bm{p}_{d,\tau}^{+}-\bm{\rho}_{-}^{\top}\bm{p}^{-}_{d,\tau}$ (2a)
s.t. $\bm{1}^{\top}(\bm{p}_{d,\tau}^{+}-\bm{p}^{-}_{d,\tau}-\bm{\kappa}_{d,\tau})=-\bm{1}^{\top}(\bm{y}_{d,\tau}-\bm{w}^{*}_{d,\tau})$ (2b)
$-\overline{\bm{f}}-\bm{H}(\bm{p}_{d,\tau}^{*}+\bm{w}^{*}_{d,\tau}-\bm{l}_{d,\tau})\leq\bm{H}(\bm{p}_{d,\tau}^{+}-\bm{p}_{d,\tau}^{-}-\bm{\kappa}_{d,\tau}$
$\qquad+\bm{y}_{d,\tau}-\bm{w}^{*}_{d,\tau})\leq\overline{\bm{f}}-\bm{H}(\bm{p}_{d,\tau}^{*}+\bm{w}^{*}_{d,\tau}-\bm{l}_{d,\tau})$ (2c)
$\bm{0}\leq\bm{p}_{d,\tau}^{+}\leq\overline{\bm{p}^{+}}$ (2d)
$\bm{0}\leq\bm{p}_{d,\tau}^{-}\leq\overline{\bm{p}^{-}}$ (2e)
$\underline{\bm{p}}\leq\bm{p}_{d,\tau}^{*}+\bm{p}_{d,\tau}^{+}-\bm{p}_{d,\tau}^{-}\leq\overline{\bm{p}}$ (2f)
$\bm{0}\leq\bm{\kappa}_{d,\tau}\leq\bm{y}_{d,\tau}$ (2g)
where $\bm{z}_{d,\tau}=[\bm{p}_{d,\tau}^{+}\in\mathbb{R}^{N};\bm{p}_{d,\tau}^{-}\in\mathbb{R}^{N};\bm{\kappa}_{d,\tau}\in\mathbb{R}^{N}]$ is the collection of RT decision variables. The output power of generators may be increased by an amount $\bm{p}_{d,\tau}^{+}$ with the marginal cost $\bm{\rho}_{+}>0$ for up-regulation, or decreased by an amount $\bm{p}_{d,\tau}^{-}$ with the marginal utility $\bm{\rho}_{-}>0$ for down-regulation. $\bm{\kappa}_{d,\tau}$ is the amount of RES spill. These decisions are driven by the need to settle the RES deviation $\bm{y}_{d,\tau}-\bm{w}_{d,\tau}^{*}$ in (2b). (2c) is the power flow constraint, whose lower and upper bounds are determined by subtracting the power flow in the DA market from the line capacity. (2d) and (2e) limit the amount of up-regulation and down-regulation power to $\overline{\bm{p}^{+}},\overline{\bm{p}^{-}}$. For inflexible generators that cannot be dispatched in RT, the corresponding elements in the upper bounds will be zero, resulting in zero up- and down-adjustments for those generators. Additionally, the eventual generation power, considering the DA schedule $\bm{p}_{d,\tau}^{*}$ and the adjustment, should be within the output power limits, as stated in (2f). The inclusion of RES spill accounts for situations where the actual RES generation surpasses the scheduled amount in the DA schedule $\bm{w}_{d,\tau}^{*}$, and the excess cannot be entirely offset by the down-regulation power available from flexible generators. The amount of RES spill $\bm{\kappa}_{d,\tau}$ can be at most to its realization $\bm{y}_{d,\tau}$, as stated in (2g).
The RT clearing at time-slot $\tau,\forall\tau=2,...,T$ is,
[…]
$\underline{\bm{r}}\leq\bm{p}_{d,\tau}^{*}+\bm{p}_{d,\tau}^{+}-\bm{p}_{d,\tau}^{-}-$
$\qquad(\bm{p}_{d,\tau-1}^{*}+\bm{p}_{d,\tau-1}^{+*}-\bm{p}_{d,\tau-1}^{-*})\leq\overline{\bm{r}}$ (3c)
The difference between the eventual generation at time-slot $\tau-1$, denoted by $\bm{p}_{d,\tau-1}^{*}+\bm{p}_{d,\tau-1}^{+*}-\bm{p}_{d,\tau-1}^{-*}$, and the eventual generation at time-slot $\tau$ must satisfy the ramping constraints, as stated in (3c).
To obtain the unique primal and dual solutions from DA and RT clearing, we require each element in the marginal cost vectors $\bm{\rho},\bm{\rho}_{+},\bm{\rho}_{-}$ are different. After solving the DA and RT market clearing, the eventual generation of the generators is either $\bm{p}_{d,\tau}^{*}+\bm{p}_{d,\tau}^{+*}$ when the RES falls short of its scheduled production in RT, or $\bm{p}_{d,\tau}^{*}-\bm{p}_{d,\tau}^{-*}$ when the RES generates more power than the schedule. Here, we define overall generation cost or the negative social surplus in a day.
Definition 1.
We define overall generation cost in a day $d$ as,
$\sum_{\tau=1}^{T}\bm{\rho}^{\top}\bm{p}^{*}_{d,\tau}+\bm{\rho}_{+}^{\top}\bm{p}^{+*}_{d,\tau}-\bm{\rho}_{-}^{\top}\bm{p}^{-*}_{d,\tau}$ (4a)
$=$ $\sum_{\tau=1}^{T}\bm{\rho}^{\top}(\bm{p}^{*}_{d,\tau}+\bm{p}^{+*}_{d,\tau})+(\bm{\rho}_{+}-\bm{\rho})^{\top}\bm{p}^{+*}_{d,\tau}+$
$\sum_{\tau=1}^{T}\bm{\rho}^{\top}(\bm{p}^{*}_{d,\tau}-\bm{p}^{-*}_{d,\tau})+(\bm{\rho}-\bm{\rho}_{-})^{\top}\bm{p}^{-*}_{d,\tau}$ (4b)
$\bm{\rho}_{+}-\bm{\rho}$ and $\bm{\rho}-\bm{\rho}_{-}$ are the incremental bidding price, which reflects the marginal opportunity loss for up- and down-regulation [5]. We require them to be positive for penalizing the RT adjustment due to forecasting errors. In this way, any RT adjustment would bring the extra cost either $(\bm{\rho}_{+}-\bm{\rho})^{\top}\bm{p}^{+*}_{d,\tau}$ or $(\bm{\rho}-\bm{\rho}_{-})^{\top}\bm{p}^{-*}_{d,\tau}$. The incremental bidding prices of supplying upward and downward balancing power are usually different. This explains why forecasting the expectation hardly works well in reducing the overall cost, as it overlooks the typical asymmetry affecting the RT cost.
II-B Mathematical Reformulation
In this subsection, we first convert the RT clearing in (2) and (3) into a mathematically equivalent form, and then give the compact form of DA and RT market clearing.
We reformulate the RT clearing in (2). To show the upper and lower bounds of the power adjustment more clearly, we divide the constraint (2f) into two parts. Concretely, when the RES produces less power than the schedule, we have $\bm{p}_{d,\tau}^{+}\geq\bm{0},\bm{p}_{d,\tau}^{-}=\bm{0}$. Conversely, when the RES produces more power than the schedule, we have $\bm{p}_{d,\tau}^{-}\geq\bm{0},\bm{p}_{d,\tau}^{+}=\bm{0}$. We divide (2f) into the following two constraints by the two cases,
$\underline{\bm{p}}-\bm{p}_{d,\tau}^{*}\leq\bm{p}_{d,\tau}^{+}\leq\overline{\bm{p}}-\bm{p}_{d,\tau}^{*}$ (5a)
$\bm{p}_{d,\tau}^{*}-\overline{\bm{p}}\leq\bm{p}_{d,\tau}^{-}\leq\bm{p}_{d,\tau}^{*}-\underline{\bm{p}}$ (5b)
Since $\underline{\bm{p}}\leq\bm{p}_{d,\tau}^{*}\leq\overline{\bm{p}}$, the left side of (5) is less than 0. Considering the power adjustment $\bm{p}_{d,\tau}^{+},\bm{p}_{d,\tau}^{-}$ is larger than 0 as stated in (2d) and (2e), (5) can be further simplified as,
$\bm{0}\leq\bm{p}_{d,\tau}^{+}\leq\overline{\bm{p}}-\bm{p}_{d,\tau}^{*}$ (6a)
$\bm{0}\leq\bm{p}_{d,\tau}^{-}\leq\bm{p}_{d,\tau}^{*}-\underline{\bm{p}}$ (6b)
The RT market clearing at time $\tau=1$ becomes,
[…]
Likewise, (3c) can be equivalently reformulated as,
$\bm{0}\leq\bm{p}_{d,\tau}^{+}\leq\overline{\bm{r}}+(\bm{p}_{d,\tau-1}^{*}+\bm{p}_{d,\tau-1}^{+*}-\bm{p}_{d,\tau-1}^{-*})-\bm{p}_{d,\tau}^{*}$ (8a)
$\bm{0}\leq\bm{p}_{d,\tau}^{-}\leq\bm{p}_{d,\tau}^{*}-(\bm{p}_{d,\tau-1}^{*}+\bm{p}_{d,\tau-1}^{+*}-\bm{p}_{d,\tau-1}^{-*})-\overline{\bm{r}}$ (8b)
The RT market clearing at time $\tau=2,...,T$ becomes,
[…]
Next, we convert the DA market clearing in (1) and the RT market clearing in (7) and (9), which are linear programs, into equivalent compact forms, with the dual variable listed after the colon. Concretely, the compact DA market clearing is,
$\bm{x}_{d}^{*}=$ $\mathop{\arg\min}_{\bm{x}_{d}}$ $\quad\bm{\rho}_{\text{DA}}^{\top}\bm{x}_{d}$ (10a)
s.t. $\quad\bm{G}_{\text{DA}}\bm{x}_{d}\leq\bm{\psi}_{\text{DA},d}+\bm{F}_{\text{DA}}^{y}\hat{\bm{y}}_{d}:\bm{\sigma}_{d}.$ (10b)
The coefficients $\bm{\rho}_{\text{DA}},\bm{G}_{\text{DA}},\bm{\psi}_{\text{DA},d},\bm{F}_{\text{DA}}^{y}$ are constant. The RES forecasts and the demand are summarized into vectors $\hat{\bm{y}}_{d}=[\hat{\bm{y}}_{d,\tau}]_{\tau=1}^{T}$ and $\bm{l}_{d}=[\bm{l}_{d,\tau}]_{\tau=1}^{T}$, respectively. The value of $\bm{\psi}_{\text{DA},d}$ varies from day to day due to its dependence on the demand $\bm{l}_{d}$. Likewise, the RT market clearing in (7) and (9) are converted into the compact forms,
$\bm{z}_{d,\tau}^{*}=$ $\mathop{\arg\min}_{\bm{z}_{d,\tau}}$ $\quad\bm{\rho}_{\text{RT}}^{\top}\bm{z}_{d,\tau}$ (11a)
s.t. $\bm{G}_{\text{RT}}\bm{z}_{d,\tau}\leq\bm{\psi}_{\text{RT},d,\tau}+\bm{F}_{\text{RT}}\bm{x}_{d,\tau}^{*}:\bm{\nu}_{d,\tau},\tau=1.$ (11b)
$\bm{z}_{d,\tau}^{*}=$ $\mathop{\arg\min}_{\bm{z}_{d,\tau}}$ $\quad\bm{\rho}_{\text{RT}}^{\top}\bm{z}_{d,\tau}$ (12a)
s.t. $\quad\bm{G}_{\text{RT}}^{\prime}\bm{z}_{d,\tau}\leq\bm{\psi}_{\text{RT},d,\tau}^{\prime}+[\bm{F}_{\text{RT}}^{\prime x},\bm{F}_{\text{RT}}^{\prime p},\bm{F}_{\text{RT}}^{\prime+-}]$
$[\bm{x}^{*}_{d,\tau},\bm{p}^{*}_{d,\tau-1},\bm{p}^{+-*}_{d,\tau-1}]^{\top}:\bm{\zeta}_{d,\tau},\forall\tau=2,...,T.$ (12b)
The coefficients $\bm{\rho}_{\text{RT}},\bm{G}_{\text{RT}},\bm{\psi}_{\text{RT},d,\tau},\bm{F}_{\text{RT}}$ and $\bm{G}^{\prime}_{\text{RT}},\bm{\psi}^{\prime}_{\text{RT},d,\tau},\bm{F}^{\prime x}_{\text{RT}},\bm{F}^{\prime p}_{\text{RT}},\bm{F}^{\prime+-}_{\text{RT}}$ are constant. The values of $\bm{\psi}_{\text{RT},d,\tau},\bm{\psi}_{\text{RT},d,\tau}^{\prime}$ vary from hour to hour due to the dependence on the RES realization $\bm{y}_{d,\tau}$. The parameter $\bm{p}^{+-*}_{d,\tau-1}=[\bm{p}^{+*}_{d,\tau-1};\bm{p}^{-*}_{d,\tau-1}]$ contains RT power adjustment for up- and down- regulation at previous time $\tau-1$.
III Methodology: Parameter Estimation
Figure 2: An illustration of the bilevel program.
In this section, we show how to identify the RIEQ, whose value corresponds to $\hat{\bm{y}}_{d,\tau}$, an input parameter in (1), via training forecasting models. At the training phase, a bilevel program [18] is formulated for estimating the forecasting model parameter. Let $g(\ \cdot\ ;\Theta)$ denote the forecasting model with the parameter $\Theta$, and $\bm{s}_{d,\tau}$ denote the context. The RIEQ forecast for the time $\tau$ on day $d$ is,
$\hat{\bm{y}}_{d,\tau}=g(\bm{s}_{d,\tau};\Theta)$ (13)
where $\hat{\bm{y}}_{d,\tau}$ denotes the forecast at the training phase, with $\Theta$ given by any value. Data in the training set $\{\{\bm{s}_{d,\tau},y_{d,\tau}\}_{\tau=1}^{T}\}_{d=1}^{D}$ is available, which is consisted of historical context and RES realization in $D$ days. An illustration of the bilevel program is shown in Fig. 2. The upper level determines the model parameter $\Theta$, while the lower level involves the DA and RT market clearings. The bilevel program is mathematically formulated as,
$\underset{{\Theta}}{\mathop{\min}}$ $\frac{1}{D\cdot T}\sum_{d=1}^{D}\{\bm{\rho}_{\text{DA}}^{\top}\bm{x}_{d}^{*}+\sum_{\tau=1}^{T}\bm{\rho}_{\text{RT}}^{\top}\bm{z}_{d,\tau}^{*}\}$ (14a)
s.t. $\hat{\bm{y}}_{d,\tau}=g(\bm{s}_{d,\tau};\Theta),\forall\tau=1,...,T,\forall d=1,...,D$ (14b)
$0\leq\hat{\bm{y}}_{d,\tau}\leq\bar{\bm{y}}_{d,\tau},\forall\tau=1,...,T,\forall d=1,...,D$ (14c)
[…]
where the upper-level objective (14a) seeks to minimize the expected overall operating cost of the two markets. This is achieved by leveraging the optimal DA and RT cost functions, which are informed by the decisions obtained from the lower level (14d). (14c) limits the forecast $\hat{\bm{y}}_{d,\tau}$ within $\bar{\bm{y}}_{d,\tau}$, which can be RES capacity. The lower level treats the forecast $\hat{\bm{y}}_{d,\tau}$ as an input parameter. As a consequence, both DA and RT decisions are affected by it.
To show the impact of the forecast on the operating cost more clearly, we replace the lower level with the dual problems. The overall operating cost within the upper-level objective is then substituted with the DA and RT dual objectives. These objectives are constructed as a linear combination of the right-side parameters and the associated dual variables, i.e,
$\underset{{\Theta}}{\mathop{\min}}$ $\frac{1}{D\cdot T}\sum_{d=1}^{D}\{-\bm{\sigma}_{d}^{*\top}(\bm{\psi}_{\text{DA},d}+\bm{F}_{\text{DA}}^{y}\hat{\bm{y}}_{d})$
$-\bm{\nu}_{d,1}^{*\top}(\bm{\psi}_{\text{RT},d,1}+\bm{F}_{\text{RT}}\bm{x}_{d,1}^{*})+$
$\sum_{\tau=2}^{T}-\bm{\zeta}_{d,\tau}^{*\top}(\bm{\psi}_{\text{RT},d,\tau}^{\prime}+\bm{F}_{\text{RT}}^{\prime x}\bm{x}_{d,\tau}^{*}+\bm{F}_{\text{RT}}^{\prime p}\bm{p}_{d,\tau-1}^{*}+$
$\bm{F}_{\text{RT}}^{\prime+-}\bm{p}_{d,\tau-1}^{+-*})\}$ (15a)
s.t. $\hat{\bm{y}}_{d,\tau}=g(\bm{s}_{d,\tau};\Theta),\forall\tau=1,...,T,\forall d=1,...,D$ (15b)
$0\leq\hat{\bm{y}}_{d,\tau}\leq\bar{\bm{y}}_{d,\tau},\forall\tau=1,...,T,\forall d=1,...,D$ (15c)
$\bm{\sigma}_{d}^{*}=\mathop{\arg\max}_{\bm{\sigma}_{d}\geq 0}-\bm{\sigma}_{d}^{\top}(\bm{\psi}_{\text{DA},d}+\bm{F}_{\text{DA}}^{y}\hat{\bm{y}}_{d}),\forall d=1,...,D$ (15d)
(10)$,\forall d=1,...,D$ (15e)
$\bm{\nu}_{d,1}^{*}=\mathop{\arg\max}_{\nu_{d,1}\geq 0}-\bm{\nu}_{d,1}^{\top}(\bm{\psi}_{\text{RT},d,1}+\bm{F}_{\text{RT}}\bm{x}_{d,1}^{*}),\forall d=1,...,D$ (15f)
(11)$,\forall d=1,...,D$ (15g)
$\bm{\zeta}_{d,\tau}^{*}=\mathop{\arg\max}_{\bm{\zeta}_{d,\tau}\geq 0}-\bm{\zeta}_{d,\tau}^{\top}(\bm{\psi}_{\text{RT},d,\tau}^{\prime}+\bm{F}_{\text{RT}}^{\prime x}\bm{x}_{d,\tau}^{*}+\bm{F}_{\text{RT}}^{\prime p}\bm{p}_{d,\tau-1}^{*}+$
$\bm{F}_{\text{RT}}^{\prime+-}\bm{p}_{d,\tau-1}^{+-*}),\forall\tau=2,...,T,\forall d=1,...,D$ (15h)
(12)$,\forall d=1,...,D$ (15i)
Since RT clearing requires the primal solutions of DA clearing and the previous RT clearing as input parameters, we also include the primal problems in the lower level. The forecast $\hat{\bm{y}}_{d}$ affects the upper-level objective (15a) via its impact on the DA and RT dual solutions $\bm{\sigma}_{d}^{*},\bm{\nu}_{d,1}^{*},\bm{\zeta}_{d,\tau}^{*}$, and their primal solutions $\bm{x}^{*}_{d,\tau},\bm{p}^{*}_{d,\tau-1},\bm{p}^{+-*}_{d,\tau-1}$. If we can obtain the function between them and the forecast $\hat{\bm{y}}_{d}$ directly, the upper-level objective can be rewritten as a function regarding the forecast $\hat{\bm{y}}_{d}$, and can be used as the loss function for training. In the next section, we will show how to achieve this.
At the operation phase, the forecast $\bm{\hat{y}}_{d,\tau}$ under the context $\bm{s}_{d,\tau}$ can be obtained by the trained model. Subsequently, utilizing this forecast, the operator proceeds to solve the DA clearing in (1), followed by the RT clearing in (2) and (3).
IV Loss Function Design
We derive the loss function based on the bilevel program in (15). We first analyze how forecasts influence the DA and RT optimal solutions. Based on it, a loss function is derived, and will be used for training a forecasting model.
IV-A The Impact of the Forecast on the DA and RT Solutions
This section analyzes the impact of the forecasts on the primal and dual solutions of DA and RT market clearings. Specifically, we derive analytical functions that quantitatively depict the impact of the forecasts on the optimal primal and dual solutions. As a parameter to the DA market clearing (10), forecasts $\hat{\bm{y}}_{d}$ influence DA primal and dual solutions directly. As for the RT market clearing, the impact of the forecast is indirect, as it does not directly appear as the parameter in (11) or (12). Concretely, for the RT market clearing at time-slot $\tau=1$, the forecast $\hat{\bm{y}}_{d}$ influences $\bm{x}_{d,1}^{*}$, and then $\bm{x}_{d,1}^{*}$ influences the RT solutions, with the first impact being determined by the DA clearing (10), and the second impact by the RT clearing (11). For the RT clearing at time-slot $\tau=2,...,T$, the impact of the forecast is more complex. The forecast $\hat{\bm{y}}_{d}$ affects the DA solutions $\bm{x}_{d,\tau}^{*},\bm{p}_{d,\tau-1}^{*}$ through the DA clearing (10). The influence of the forecast $\hat{\bm{y}}_{d}$ on the RT solutions $\bm{p}_{d,\tau-1}^{+-*}$ at the previous time-slot $\tau-1$ occurs as explained earlier. Then, the influence of the parameters $\bm{x}_{d,\tau}^{*},\bm{p}^{*}_{d,\tau-1},\bm{p}^{+-*}_{d,\tau-1}$ on the RT clearing solutions at time $\tau$ is determined by (12). The above process is summarized in Fig. 3.
Figure 3: An illustration of the impact of forecasts on DA and RT primal solutions.
Since the core of the above analysis is understanding the impact of the parameters on the optimization solutions, we use the multiparametric theory for this end. A general linear program (16) is used as an example. We firstly define primal and dual decision policies. Then, the theorem regarding them is presented.
$\bm{x}^{*}=$ $\mathop{\arg\min}_{\bm{x}}$ $\quad\bm{c}^{\top}\bm{x}$ (16a)
s.t. $\quad\bm{G}\bm{x}\leq\bm{\psi}+\bm{F}\bm{\omega}:\bm{\sigma}.$ (16b)
Definition 2.
(Primal and dual decision policies) Primal and dual decision policies are functions defined across the polyhedral set $\Omega$, which describe the change in the optimal primal and dual solutions, i.e., $\bm{x}^{*}$ and $\bm{\sigma}^{*}$, as the parameter $\bm{\omega}$ varies in $\Omega$.
Theorem 1.
[22] Consider the linear program (16) and the parameter $\bm{\omega}\in\Omega$. The primal and dual decision policies are a piecewise linear function and a stepwise function respectively, if there exists a polyhedral partition $R_{1},...,R_{N}$ of $\Omega$, and $\forall\bm{\omega}\in R_{i}$, the primal decision policy is linear, and the dual decision policy is a constant function.
Theorem 1 implies that in a neighborhood of the parameter, the primal decision policy is represented by a linear function, whereas the corresponding dual decision policy remains a constant function. Given a specific value of $\bm{w}$, we study the local policies defined in its neighborhood. The local dual decision policy can be obtained easily, as it is a constant function. Its output is the optimal dual solution obtained by solving the dual problem of (16), given the value of $\bm{\omega}$. Additionally, after solving (16), the active constraints of (16b) can be obtained. Let $\mathcal{J}$ denote the row index set associated with (16b), and $\mathcal{J}^{a}$ denote the row index set of the active constraints. The parameters associated with the active constraints are denoted as $\bm{G}[\mathcal{J}^{a}],\bm{\psi}[\mathcal{J}^{a}],\bm{F}[\mathcal{J}^{a}]$. They are the sub-matrices and sub-vectors of $\bm{G},\bm{\psi},\bm{F}$ and are comprised by the rows of $\bm{G},\bm{\psi},\bm{F}$ in the row index sets $\mathcal{J}^{a}$. We have the following proposition for the local primal decision policy,
Proposition 1.
The local primal decision policy of (16) is,
$\bm{x}^{*}=\bm{G}[\mathcal{J}^{a}]^{-1}(\bm{\psi}[\mathcal{J}^{a}]+\bm{F}[\mathcal{J}^{a}]\bm{\omega})$ (17)
Proof.
With the active constraints of (16), we have $\bm{G}[\mathcal{J}^{a}]\bm{x}^{*}=\bm{\psi}[\mathcal{J}^{a}]+\bm{F}[\mathcal{J}^{a}]\bm{\omega}$. $\bm{G}[\mathcal{J}^{a}]^{-1}$ is the pseudo inverse of $\bm{G}[\mathcal{J}^{a}]$. By moving $\bm{G}[\mathcal{J}^{a}]$ from left to right, we have (17). ∎
Remark 2.
(17) calculates the inverse of $\bm{G}[\mathcal{J}^{a}]$, whose computational complexity depends on the matrix size, i.e., $|\mathcal{J}^{a}|$. We note that a similar matrix inverse is also involved in [14], with the matrix size of $|\mathcal{J}|+\iota$, where $\iota$ is the dimension of $\bm{x}$. There is $|\mathcal{J}|+\iota>|\mathcal{J}^{a}|$. Therefore, the computation burden of [14] is larger.
With Proposition 1, we present the local primal decision policy for the DA clearing (10), and RT clearing (11) and (12). Let $\mathcal{J}^{a}_{\text{DA},d},\mathcal{J}^{a}_{\text{RT},d,\tau},\forall\tau=1,...,T$ denote the row index set of active constraints (10b), (11b), and (12b). The parameters associated with the active constraints are denoted as $\bm{G}_{\text{DA}}[\mathcal{J}_{\text{DA},d}^{a}],\bm{\psi}_{\text{DA},d}[\mathcal{J}_{\text{DA},d}^{a}],\bm{F}_{\text{DA}}^{y}[\mathcal{J}_{\text{DA},d}^{a}]$, $\bm{G}_{\text{RT}}[\mathcal{J}_{\text{RT},d,\tau}^{a}],\bm{\psi}_{\text{RT},d,\tau}[\mathcal{J}_{\text{RT},d,\tau}^{a}],\bm{F}_{\text{RT}}[\mathcal{J}_{\text{RT},d,\tau}^{a}]$,$\bm{G}_{\text{RT}}^{\prime}[\mathcal{J}_{\text{RT},d,\tau}^{a}]$, and $\bm{\psi}_{\text{RT},d,\tau}^{\prime}[\mathcal{J}_{\text{RT},d,\tau}^{a}],\bm{F}_{\text{RT}}^{\prime x}[\mathcal{J}_{\text{RT},d,\tau}^{a}],\bm{F}_{\text{RT}}^{\prime p}[\mathcal{J}_{\text{RT},d,\tau}^{a}],\bm{F}_{\text{RT}}^{\prime+-}[\mathcal{J}_{\text{RT},d,\tau}^{a}]$. We have the following proposition for the local primal decision policies of (10), (11), and (12).
Proposition 2.
The local primal decision policies of (10),(11) and (12) are,
$f^{x}_{d}(\hat{\bm{y}}_{d}):=\bm{x}_{d}^{*}=\bm{G}_{\text{DA}}[\mathcal{J}_{\text{DA},d}^{a}]^{-1}(\bm{\psi}_{\text{DA},d}[\mathcal{J}_{\text{DA},d}^{a}]+\bm{F}_{\text{DA}}^{y}[\mathcal{J}_{\text{DA},d}^{a}]\hat{\bm{y}}_{d})$ (18)
$\begin{split}&\bm{z}_{d,\tau}^{*}=\bm{G}_{\text{RT}}[\mathcal{J}_{\text{RT},d,\tau}^{a}]^{-1}(\bm{\psi}_{\text{RT},d,\tau}[\mathcal{J}_{\text{RT},d,\tau}^{a}]+\\ &\bm{F}_{\text{RT}}[\mathcal{J}^{a}_{\text{RT},d,\tau}]\bm{x}_{d,\tau}^{*}),\tau=1\end{split}$ (19)
$\begin{split}&\bm{z}_{d,\tau}^{*}=\bm{G}_{\text{RT}}^{\prime}[\mathcal{J}_{\text{RT},d,\tau}^{a}]^{-1}(\bm{\psi}^{\prime}_{\text{RT},d,\tau}[\mathcal{J}_{\text{RT},d,\tau}^{a}]+\bm{F}_{\text{RT}}^{\prime x}[\mathcal{J}_{\text{RT},d,\tau}^{a}]\bm{x}_{d,\tau}^{*}+\\ &\bm{F}_{\text{RT}}^{\prime p}[\mathcal{J}_{\text{RT},d,\tau}^{a}]\bm{p}_{d,\tau-1}^{*}+\bm{F}_{\text{RT}}^{\prime+-}[\mathcal{J}_{\text{RT},d,\tau}^{a}]\bm{p}_{d,\tau-1}^{+-*}),\forall\tau=2,...,T\end{split}$ (20)
Eq. (18) is a linear function of $\hat{\bm{y}}_{d}$, whose output is the DA solution $\bm{x}_{d}^{*}$ on the day $d$. In the following, we show how to rewrite (19) and (20) as a function of the forecast $\hat{\bm{y}}_{d}$. The illustration is in Fig. 4.
Figure 4: An illustration of the calculation for rewriting (19) for $\tau=1$ and (20) for $\tau=2$ as functions of the forecast $\hat{\bm{y}}_{d}$. Different functions are represented by boxes in various colors. The diamond represents the calculation of the operator $\Pi_{\mathcal{J}}$ to obtain the function relating part of the vector to the forecast $\bm{\hat{y}}_{d}$. The circle represents the linear transformation.
In addition, we define an operator $\Pi_{\mathcal{J}}:h(\bm{x})\mapsto\tilde{h}(\bm{x})$, where $h(\bm{x})=\bm{A}\bm{x}+\bm{b}$ is a linear function with the parameter $\bm{A},\bm{b}$, and $\mathcal{J}$ is the row index subset of $\bm{A},\bm{b}$. The output $\tilde{h}(\bm{x})$ of the operator is also a linear function, where $\tilde{h}(\bm{x})=\bm{A}[\mathcal{J}]\bm{x}+\bm{b}[\mathcal{J}]$ and $\bm{A}[\mathcal{J}],\bm{b}[\mathcal{J}]$ are the sub-matrix and sub-vector of $\bm{A},\bm{b}$.
Let $\mathcal{I}_{\text{DA},d,\tau}$ denote the row index set corresponding to $\bm{x}_{d,\tau}^{*}$ within $\bm{x}_{d}^{*}$. The function that maps $\hat{\bm{y}}_{d}$ to $\bm{x}_{d,\tau}^{*}$ is,
$\begin{split}&f^{x}_{d,\tau}(\hat{\bm{y}}_{d}):=\bm{x}_{d,\tau}^{*}=\Pi_{\mathcal{I}_{\text{DA},d,\tau}}(f^{x}_{d}(\hat{\bm{y}}_{d})),\forall\tau=1,...,T\end{split}$ (21)
The coefficients are determined by the coefficients of $f^{x}_{d}(\hat{\bm{y}}_{d})$ in (18), whose row indexes belong to the set $\mathcal{I}_{\text{DA},d,\tau}$. By substituting (21) into (19), we can obtain the function between RT primal solution $\bm{z}_{d,\tau}^{*}$ at time $\tau=1$ and the forecast $\hat{\bm{y}}_{d}$.
$\begin{split}&f_{d,\tau}^{z}(\hat{\bm{y}}_{d}):=\bm{z}_{d,\tau}^{*}=\bm{G}_{\text{RT}}[\mathcal{J}_{\text{RT},d,\tau}^{a}]^{-1}\bm{F}_{\text{RT}}[\mathcal{J}^{a}_{\text{RT},d,\tau}]f_{d,\tau}^{x}(\hat{\bm{y}}_{d})+\\ &\bm{G}_{\text{RT}}[\mathcal{J}_{\text{RT},d,\tau}^{a}]^{-1}\bm{\psi}_{\text{RT},d,\tau}[\mathcal{J}_{\text{RT},d,\tau}^{a}],\tau=1\end{split}$ (22)
To obtain the function between RT primal solution $\bm{z}_{d,\tau}^{*}$ at time-slot $\tau=2,...,T$ and the forecast $\hat{\bm{y}}_{d}$, we need to obtain the function between DA solution $\bm{p}_{d,\tau-1}^{*}$, RT solution $\bm{p}_{d,\tau-1}^{+-*}$ and the forecast $\hat{\bm{y}}_{d}$ as well. Concretely, since $\bm{p}_{d,\tau-1}^{*}$ is a part of $\bm{x}_{d,\tau-1}^{*}$, let $\mathcal{I}_{\text{DA},d,\tau-1}^{p}$ be the row index set corresponds to $\bm{p}_{d,\tau-1}^{*}$ within $\bm{x}_{d,\tau-1}^{*}$. With (21), the function between $\bm{p}_{d,\tau-1}^{*}$ and the forecast $\hat{\bm{y}}_{d}$ is,
$\begin{split}&f^{p}_{d,\tau-1}(\hat{\bm{y}}_{d}):=\bm{p}_{d,\tau-1}^{*}=\Pi_{\mathcal{I}_{\text{DA},d,\tau-1}^{p}}(f_{d,\tau-1}^{x}(\hat{\bm{y}}_{d})),\\ &\forall\tau=2,...,T\end{split}$ (23)
Likewise, since $\bm{p}_{d,\tau-1}^{+-*}$ is a part of $\bm{z}_{d,\tau-1}^{*}$, let $\mathcal{I}_{\text{RT},d,\tau-1}^{+-}$ be the row index set corresponds to $\bm{p}_{d,\tau-1}^{+-*}$ within $\bm{z}_{d,\tau-1}^{*}$. We can express the function of $\bm{p}_{d,\tau-1}^{+-*}$ w.r.t. $\hat{\bm{y}}_{d}$ as,
$\begin{split}&f_{d,\tau-1}^{+-}(\hat{\bm{y}}_{d}):=\bm{p}_{d,\tau-1}^{+-*}=\Pi_{\mathcal{I}_{\text{RT},d,\tau-1}^{+-}}(f_{d,\tau-1}^{z}(\hat{\bm{y}}_{d})),\\ &\forall\tau=2,...,T\end{split}$ (24)
Accordingly, by substituting (21), (23), (24) into (20), the function between $\bm{z}_{d,\tau}^{*}$ at time-slot $\tau=2,...,T$ is,
$\begin{split}&f^{z}_{d,\tau}(\hat{\bm{y}}_{d}):=\\ &\bm{z}_{d,\tau}^{*}=\bm{G}_{\text{RT}}^{\prime}[\mathcal{J}_{\text{RT},d,\tau}^{a}]^{-1}(\bm{\psi}^{\prime}_{\text{RT},d,\tau}[\mathcal{J}_{\text{RT},d,\tau}^{a}]+\\ &\bm{F}_{\text{RT}}^{\prime x}[\mathcal{J}_{\text{RT},d,\tau}^{a}]f_{d,\tau}^{x}(\hat{\bm{y}}_{d})+\bm{F}_{\text{RT}}^{\prime p}[\mathcal{J}_{\text{RT},d,\tau}^{a}]f_{d,\tau-1}^{p}(\hat{\bm{y}}_{d})+\\ &\bm{F}_{\text{RT}}^{\prime+-}[\mathcal{J}_{\text{RT},d,\tau}^{a}]f_{d,\tau-1}^{+-}(\hat{\bm{y}}_{d})),\forall\tau=2,...,T\end{split}$ (25)
To sum up, the linear functions between the forecast $\hat{\bm{y}}_{d}$ and DA and RT primal solutions, which are defined in the neighborhood of $\hat{\bm{y}}_{d}$, are summarized in the (21)-(25). The function between the optimal dual solutions and the forecast $\hat{\bm{y}}_{d}$ is a constant function, whose output can be obtained by solving the dual problems of (10),(11), and (12). With these, we are ready to transform the upper-level objective (15a) to a function of the forecast $\hat{\bm{y}}_{d}$. The details are in the next subsection.
IV-B Loss Function Derivation
By substituting the functions (21),(23),(24) and dual solutions into the upper-level objective (15a), the loss function in the neighborhood of the forecast $\hat{\bm{y}}_{d}$ is,
$\ell_{d}(\hat{\bm{y}}_{d}):=-\bm{\sigma}_{d}^{*\top}(\bm{\psi}_{\text{DA},d}+\bm{F}_{\text{DA}}^{y}\hat{\bm{y}}_{d})-\bm{\nu}_{d,1}^{*\top}(\bm{\psi}_{\text{RT},d,1}+$
$\bm{F}_{\text{RT}}f_{d,1}^{x}(\hat{\bm{y}}_{d}))+\sum_{\tau=2}^{T}-\bm{\zeta}_{d,\tau}^{*\top}(\bm{\psi}_{\text{RT},d,\tau}^{\prime}+\bm{F}_{\text{RT}}^{\prime x}f_{d,\tau}^{x}(\hat{\bm{y}}_{d})+$
$\bm{F}_{\text{RT}}^{\prime p}f_{d,\tau-1}^{p}(\hat{\bm{y}}_{d})+\bm{F}_{\text{RT}}^{\prime+-}f_{d,\tau-1}^{+-}(\hat{\bm{y}}_{d}))$ (26a)
[…]
Since functions in (26b) are linear, and the dual decision policies are constant, the loss function $\ell_{d}(\hat{\bm{y}}_{d})$ in the neighborhood of the forecast $\hat{\bm{y}}_{d}$ is linear, and outputs the overall operating cost (4a) given the forecast $\hat{\bm{y}}_{d}$. Naturally, the derivative of $\ell_{d}(\hat{\bm{y}}_{d})$ w.r.t. the forecast $\hat{\bm{y}}_{d}$, i.e., $\frac{\partial\ell_{d}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}$, measures the marginal impact of the forecast on the overall cost and is a constant. It is,
$\begin{split}&(\frac{\partial\ell_{d}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}})^{\top}=-\bm{\sigma}_{d}^{*\top}\bm{F}_{\text{DA}}^{y}-\bm{\nu}_{d,1}^{*\top}\bm{F}_{\text{RT}}\frac{\partial f_{d,1}^{x}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}+\sum_{\tau=2}^{T}-\bm{\zeta}_{d,\tau}^{*T}\\ &(\bm{F}_{\text{RT}}^{\prime x}\frac{\partial f_{d,\tau}^{x}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}+\bm{F}_{\text{RT}}^{\prime p}\frac{\partial f_{d,\tau-1}^{p}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}+\bm{F}_{\text{RT}}^{\prime+-}\frac{\partial f_{d,\tau-1}^{+-}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}})\end{split}$ (27)
where the derivatives $\frac{\partial f_{d,\tau}^{x}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}$, $\frac{\partial f_{d,\tau-1}^{p}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}$, $\frac{\partial f_{d,\tau-1}^{+-}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}$ in (27) are constants, which measure the marginal impact of the forecast on the primal solutions. We have the following proposition regarding the loss function.
Proposition 3.
The loss function defined across the entire space of $\hat{\bm{y}}_{d}$ is a piecewise linear function. Each piece is a linear function described in (26).
Proof.
According to Theorem 1, (21),(23),(24) are piecewise linear functions defined across the entire space of $\hat{\bm{y}}_{d}$. Since the loss function is the linear transformation of (21),(23),(24), it is also a piecewise linear function. ∎
Since (21),(23),(24) are determined by different active constraint index sets $\mathcal{J}^{a}_{\text{DA},d},\mathcal{J}^{a}_{\text{RT},d,\tau-1},\forall\tau=2,...,T$ of DA and RT market clearing, each piece of the loss function is related with different active constraint index sets we well. One way is to enumerate all possible active constraint index sets, and derive the corresponding piece of the loss function [23]. However, the enumeration can be computationally expensive, particularly when dealing with large-scale optimization.
Therefore, we propose to derive the pieces of the loss function in an on-demand manner. Concretely, during training, we only recalculate the piece of the loss function when encountering new active index sets. The details are given in the next Section.
V Solution Strategy
We illustrate the training phase of the forecasting model based on neural networks (NNs). With the loss function, we use batch optimization to train NN. Given a batch of data in $B$ days, the parameter estimation with the derived loss function is formulated as,
$\underset{{\Theta}}{\mathop{\min}}$ $\frac{1}{B\cdot T}\sum_{d=1}^{B}\ell_{d}(\hat{\bm{y}}_{d})$ (28a)
s.t. $\hat{\bm{y}}_{d,\tau}=g(\bm{s}_{d,\tau};\Theta),\forall\tau=1,...,T,\forall d=1,...,B$ (28b)
$0\leq\hat{\bm{y}}_{d,\tau}\leq\bar{\bm{y}}_{d,\tau},\forall\tau=1,...,T,\forall d=1,...,B$ (28c)
Different from the conventional unconstrained program at the training phase, (28) is with the box constraint (28c) for the NN output. We design a specific model structure to address this. Specifically, a Sigmoid function, whose output is between 0 and 1, is used as the activation function at the output layer. By multiplying its output with the cap $\bar{\bm{y}}_{d,\tau}$ of each sample, the constraint (28c) is satisfied. NN’s structure is in Fig. 5.
Figure 5: The structure of neural network.
During the training, the NN, with the parameter $\Theta$ given by any value, outputs the forecast $\hat{\bm{y}}_{d,\tau}$ by (13). The primal and dual problems of DA and RT clearing in (10),(11),(12) are solved. The active constraint index sets $\mathcal{J}^{a}_{\text{DA},d},\mathcal{J}^{a}_{\text{RT},d,\tau-1},\forall\tau=2,...,T$ and the dual solutions $\bm{\sigma}_{d}^{*},\bm{\nu}_{d,1}^{*},\bm{\zeta}_{d,\tau}^{*},\forall\tau=2,...,T$ are obtained. We check if the active constraint index sets are in the buffers $\Lambda_{\text{DA}}\Lambda_{\text{RT},\tau-1},\forall\tau=2,...,T$. If yes, the stored derivatives $\frac{\partial f_{d,\tau}^{x}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}$, $\frac{\partial f_{d,\tau-1}^{p}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}$, $\frac{\partial f_{d,\tau-1}^{+-}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}$ in the buffers are used for calculating the gradient $\frac{\partial\ell_{d}(\hat{\bm{y}}_{d})}{\partial\hat{\bm{y}}_{d}}$ in (27). If not, we calculate the derivatives associated with the new ones, and calculate the gradient in (27). Additionally, we store the new active index sets and the associated derivatives in the buffers. The illustration of gradient calculation is in Fig. 6.
Figure 6: The illustration of gradient calculation in the training process.
VI Case Study
[…]
Figure 7: The illustration of IEEE 9-bus system.
Table I: Cost and technical data in DA and RT market of IEEE 9-Bus system.
$G_{1}$ $G_{2}$ $G_{3}$
Marginal generation cost $\bm{\rho}$ (\$/MW) 20 22 24
Minimum generation power $\underline{\bm{p}}$ (MW) 0 0 0
DA market Maximum generation power $\overline{\bm{p}}$ (MW) 150 200 270
Lower ramping limits $\underline{\bm{r}}$ (MW) -90 -80 -70
Upper ramping limits $\overline{\bm{r}}$ (MW) 90 80 70
Marginal cost $\bm{\rho}_{+}$ (\$/MW) 50 52 54
Marginal utility $\bm{\rho}_{-}$ (\$/MW) 18 16 14
Marginal opportunity loss for up-regulation $\bm{\rho}_{+}-\bm{\rho}$ (\$/MW) 30 30 30
RT Market Marginal opportunity loss for down-regulation $\bm{\rho}-\bm{\rho}_{-}$ (\$/MW) 2 6 10
Up-regulation limit $\overline{\bm{p}^{+}}$ (MW) 60 60 60
Down-regulation limit $\overline{\bm{p}^{-}}$ (MW) 60 60 60
We use a four-layer ResNet as the forecasting model, which has 256 hidden layer units. Its structure is described in Figure 5. The input context consists of the numeric weather prediction (i.e., the predicted wind speed and direction at 10m and 100m altitude) of each wind farm in the system. We use Root Mean Squared Error (RMSE) on the test set for assessing the forecast quality, and the average overall operating cost on the test set, as defined in (4a), for evaluating the operation value. Four benchmark models are used for comparison: Two quality-oriented (Qua-E and Qua-Q), one value-oriented point forecasting approach (OptNet), and a stochastic market clearing (Sto-OPT-P), i.e.,
  1. 1.
    Qua-E: The forecasting model (ResNet) is trained with the loss function of Mean Squared Error (MSE). The trained model predicts conditional expectations.
  2. 2.
    Qua-Q: The forecasting model is trained with the pinball loss. The trained model offers quantile predictions. Light Gradient Boosted Machine, the winner of GEFCom 2014 [25], is used as the forecasting model.
  3. 3.
    OptNet: The forecasting model (ResNet) is trained by differentiable programming [15, 14], to minimize overall system operating costs.
  4. 4.
    Sto-OPT-P: For stochastic clearing, 50 wind power scenarios are obtained via k-nearest-neighbors, which is a Prescriptive approach [26]. For each sample on the test set, the stochastic clearing is for settling the schedule of generators and wind power in DA. Then, the adjustment in (11) and (12) are performed in RT.
VI-A The Operational Advantage
[…]
Since Sto-OPT-P does not need training or relies on a point forecast, its RMSE is not reported. Sto-OPT-P serves as the ideal benchmark [7], which has the least average operating cost. The proposed approach outperforms all other methods in terms of average operating cost on the test set, except for Sto-OPT-P. However, its test time is much shorter than Sto-OPT-P, demonstrating computational efficiency. Also, we observe that the performance of Sto-OPT-P is heavily influenced by the number of scenarios used. When fewer scenarios, such as 20, are employed, the average operating cost on the test set increases to \$84,478, which is even worse than that achieved by the proposed approach.
The proposed approach exhibits a higher RMSE compared to Qua-E. This underscores the point that more accurate forecasts don’t always translate to better operational performance. Additionally, since the incremental bidding price $\bm{\rho}_{+}-\bm{\rho}$ is larger than $\bm{\rho}-\bm{\rho}_{-}$, the marginal opportunity loss of up-regulation is larger than that of down-regulation. Therefore, Qua-Q, which issues the quantile forecasts with a low proportion level ($\frac{1}{16}$), has better performance than Qua-E. As for the training time, the proposed approach requires a longer training time than Qua-E due to the more complex computation involved in calculating the gradient during the training process. But it is still acceptable, and much shorter than the value-oriented forecasting approach OptNet.
Table II: RMSE, average operating cost on the test set and the training/test time of the proposed approach and benchmarks.
Proposed Qua-E Qua-Q OptNet Sto-OPT-P
[…]
Training time per epoch 25 s 0.08 s - 136 s -
Test time 5 s 5 s 5 s 5 s 64min
VI-B The Sensitivity Analysis
In this section, we compare the proposed approach against Qua-E under different wind power capacities and marginal cost $\bm{\rho}^{+}$ for up-regulation.
VI-B1 Performance under Different Wind Power Penetration
[…]
Figure 8: Average operating cost under different wind power capacities. Qua-E stands for quality-oriented forecast issuing expectation.
VI-B2 Performance under Different Up-regulation Cost in RT
[…]
To sum up, the operation advantage of the proposed approach is more evident, under large penetration of wind power, and high up-regulation marginal cost.
Table III: Average operating cost in different settings of up-regulation costs in RT. The first column gives the RT up-regulation costs and marginal opportunity loss in each setting. Inside the parenthesis, the up-regulation marginal costs are given on the left of the slash, and the marginal opportunity loss is given on the right.
Settings Proposed (\$) Qua-E (\$) Cost reduction (%)
High up- regulation marginal cost ($G_{1}$: 80/60, $G_{2}$: 82/60, $G_{3}$: 84/60) 85114 92486 8%
Low up- regulation marginal cost ($G_{1}$: 21/1, $G_{2}$: 23/1, $G_{3}$: 25/1) 81517 81677 0.2%
Figure 9: The 6-day wind power forecast profiles for a wind farm connected to Node 5 are provided by Qua-E and the proposed approach under both high and low up-regulation cost settings.
VI-C Extending to Larger Scale System
We test the proposed approach on the IEEE 39-bus system, with 21 loads, 2 wind farms at buses 38 and 39, and 10 generators. The illustration of the IEEE 39-bus system and parameters for generators are given in our Github repo [24]. The average operating costs of the proposed one and Qua-E are shown in Fig. 8(b). The proposed approach results in lower average operating costs than Qua-E. With increased wind power capacity in the IEEE 39-bus system, the advantage of the proposed approach in decreasing the overall costs is even more pronounced compared to the IEEE 9-bus system.
VII Conclusions
We propose a value-oriented renewable energy forecasting approach, for minimizing the expected overall operating cost in the existing deterministic market clearing framework. We analytically derive the loss function for value-oriented renewable energy forecasting in sequential market clearing. The loss function is proved to be piecewise linear when the market clearing is modeled by linear programs. Additionally, we provide the analytical gradient of the loss function with respect to the forecast, which leads to an efficient training strategy. In the case study, compared to quality-oriented forecasting approach trained by MSE, the proposed approach can reduce average operating cost on the test set to 2.9% for the IEEE 9-bus system. Such an advantage is more obvious under large wind power capacity and high up-regulation costs. Under high up-regulation marginal costs, our approach can reduce the cost by up to 8%. Future work will include derivation of the value-oriented loss function for market clearing modeled by other types of optimization programs, such as quadratic optimization and conic optimization.